# Tournament design: A review from an operational research perspective

### v1

@@ 2.3.2 Fairness
In contrast to efficacy, there is no uniform definition of fairness as there exists no universally agreed system of ethics (Palacios-Huerta,, 2023). It also depends on the design element that is considered to a great extent. Still, we can observe similarities between the different ways that fairness is defined in the literature. We have identified two main principles of fairness. The first one is based on the idea that, a priori, a player should not be disadvantaged compared to any weaker player. This means that, for every player, the probability of winning the tournament should be higher than that of any weaker player. This concept is also commonly called order preservation (Schwenk,, 2000; Groh et al.,, 2008; Vu and Shoham,, 2011; Prince et al.,, 2013; Karpov,, 2016; Placek,, 2023).
@@ 2.3.3 Attractiveness
On the other hand, the importance of a match is usually defined with respect to a certain prize. Jennett, (1984) suggested a simple measure to estimate the uncertainty of winning the championship ex post. Let $k$ be the total number of points that turned out to be necessary to win the championship. The importance of a match is the inverse of the minimum number of matches required to achieve $k$ points. This metric was later used by Borland and Lye, (1992) and Dobson and Goddard, (1992) to investigate attendance rates. A decade later, Schilling, (1994) introduced probably the most well-known importance measure. The Schilling importance ${}_{i}S(X)_{t,t+k}$ of match $t+k$ at time period $t$ for team $i$ is defined as:
@@
Many other ways have been suggested to quantify match importance. Goossens et al., (2012) take into account the number of teams that are still competing for a given prize. An integer program is used to determine the highest and lowest rank any team can still obtain at the end of the season when a certain match is played. Similar formulations are proposed by Ribeiro and Urrutia, (2005); Raack et al., (2014); Gotzes and Hoppmann, (2022). Geenens, (2014) measures the decisiveness of a match based on the expected entropy (i.e. uncertainty). Corona et al., (2017) use an entropy-based measure to identify the most decisive matches in the 2014 FIFA World Cup and the 2015 Copa América. Goller and Heiniger, (2023) extend the entropy measure by considering the difference between the reward probability distributions derived from the possible outcomes of a single event. Csató et al., (2024) introduce a classification scheme to determine weakly (where the final rank of one team is known, regardless of the match outcome) and strongly (where the match outcome influences neither of both teams) stakeless matches.
@@ 3 Tournament format
The comparison of tournament formats has received much attention in operational research since determining the format is probably the first task of the contest organizer. Even if the number of games that can be played is strictly limited, usually several options are available. For example, Guyon, (2020) uncovers a serious drawback of the original format of the 2026 FIFA World Cup and lists 7 alternative options. Rennó-Costa, (2023) presents another one based on double elimination, while Guajardo and Krumer, (2023) develop three further formats. Choosing the best from such a set of feasible designs clearly requires their evaluation with respect to a number of criteria.
3.1 Effectivity and efficacy
@@
At its inception, the tournament design literature was mainly concerned with the effectivity of tournament formats. Pioneering work was done by David, (1959), who used combinatorial techniques to compute exact winning probabilities for a SKO and SRR. To that end, a series of pairwise win probabilities needs to be assumed; $\pi_{ij}$ is the probability that $i$ beats $j$, which are collected in matrix $\Pi$. David, (1959) uses a cumulative normal distribution $\Phi(\cdot)$ for the pairwise win probabilities, i.e. $\pi_{ij}=\Phi\{\frac{1}{\sigma}(r_{i}-r_{j})\}$, with $\sigma$ being the variance term and $r_{i}$ being the “true” rating of player $i$. Glenn, (1960) and Searls, (1963) apply a similar approach to evaluate tournament structures, although they assume an arbitrary matrix $\Pi$ instead of estimating the probabilities through a distribution function. All three papers compute the probabilities that other players finish in a certain position as well. Although limited to tournaments with 8 players, the authors find SKO to be the least effective format. Double and serial knockout formats are more effective, in particular, if matches are played as a best of three. For random knockout tournaments, Knuth and Lossers, (1987) compute the chance of player $i$ winning the tournament, assuming that it wins against player $i+1$ with probability $p$ and always defeats player $j$ if $j>i+1$. Adler et al., (2017) derive a lower bound on the winning probability for the best player, as well as upper and lower bounds for all players if player $i$ wins against player $j$ with probability $v_{i}/\left(v_{i}+v_{j}\right)$, where the values $v_{1}$, …, $v_{n}$ are exogenously given.
@@
The exact computation of these expressions takes all possible pairwise match outcomes into account, which quickly becomes intractable. Therefore, instead of computing the exact probabilities, the standard approach in the literature is to perform Monte Carlo simulation. Using this approach, several papers conclude that round robin tournaments are typically more efficacious than knockout tournaments (Appleton,, 1995; McGarry and Schutz,, 1997; Ryvkin and Ortmann,, 2008; Scarf et al.,, 2009; Ryvkin,, 2010; Scarf and Yusof,, 2011; Csató, 2021b, ). McGarry and Schutz, (1997) and Stanton and Williams, (2013) reveal that the efficacy of a SKO will increase by extending it towards a DKO. Efficacy can usually be improved by increasing the number of matches (Lasek and Gagolewski,, 2018). However, according to Csató, 2021b , the relationship is not necessarily monotonic for complex hybrid tournaments. According to earlier studies (Appleton,, 1995; Dinh et al.,, 2020), the Swiss system is rather weak in identifying the true ranking of players. However, Sziklai et al., (2022) find that the Swiss system outperforms other formats with the same number of games. To the best of our knowledge, the efficacy of incomplete round robin tournaments has not been examined yet.
@@
215364
[…]
32154
Figure 2: Structures in unbalanced knockout trees (Arlegi and Dimitrov,, 2020)
[…]
Table 1: Papers on the effectivity of tournament formats
Study
Format
$\pi_{ij}$
Best Worst $n$
David, (1959)
[…]
[l]SRR, SKO,
serial-SKO
$\Phi\{\left(r_{i}-r_{j}\right)/\sigma\}$
serial-SKO
NS 3-8
Glenn, (1960)
[…]
[l]SRR, SKO, SKO-3,
serial-SKO, DKO
$\Pi^{a}$
SKO-3
SKO 4
Searls, (1963)
[…]
[l]SKO, SKO-3,
SKO+SKO,
SKO-3+SKO-3
$\Pi^{a}$
SKO-3+SKO-3
SKO 8
Appleton, (1995)
@@
SKO 8, 16
Ryvkin and Ortmann, (2008)
SKO, SRR
pdf1
SRR SKO NS
Ryvkin, (2010)
SKO, SRR
pdf1
SRR SKO NS
Table 2: Papers on the efficacy of tournament formats
@@
Unsurprisingly, there exist unavoidable trade-offs between these fairness criteria. The distance minimization and the break maximization problems are equivalent if the distance between every pair of teams is equal (Urrutia and Ribeiro,, 2006). Craig et al., (2009) focus on the trade-off between minimizing travel distance for teams, avoiding sequences of consecutive home or away games, and ensuring equity in rest times. For a SRR with $n$ teams, the minimal number of breaks is $n-2$, which is achieved (for instance) by the so-called canonical schedule. However, this schedule also yields maximally unbalanced carry-over effects for an even number of teams (Lambrechts et al.,, 2018). Günneç and Demir, (2019), Goossens et al., (2020), and Çavdaroğlu and Atan, (2022) discuss the trade-off between minimizing breaks and balancing the carry-over effect. Obviously, there are also trade-offs between fairness and organizational or financial criteria (see e.g. Goossens and Spieksma, (2009); Ribeiro and Urrutia, (2012)).

### v2

@@ 2.3.2 Fairness
In contrast to efficacy, there is no uniform definition of fairness as there exists no universally agreed system of ethics (Baumer, , 2024; Palacios-Huerta, , 2023). Still, we have identified two main principles of fairness. The first one is based on the idea that, a priori, a player should not be disadvantaged compared to any weaker player. This means that, for every player, the probability of winning the tournament should be higher than that of any weaker player. This concept is also commonly called order preservation (Schwenk, , 2000; Groh et al., , 2008; Vu and Shoham, , 2011; Prince et al., , 2013; Karpov, , 2016; Placek, , 2023).
@@ 2.3.3 Attractiveness
On the other hand, the importance of a match is usually defined with respect to a certain prize. Jennett, (1984) suggested a simple measure to estimate the uncertainty of winning the championship ex post. Let $k$ be the total number of points that turned out to be necessary to win the championship. The importance of a match is the inverse of the minimum number of matches required to achieve $k$ points. This metric was later used by Borland and Lye, (1992) and Dobson and Goddard, (1992) to investigate attendance rates. A decade later, Schilling, (1994) introduced probably the importance measure that is probably most well-known. The Schilling importance ${}_{i}S(X)_{t,t+k}$ of match $t+k$ at time period $t$ for team $i$ is defined as:
@@
Many other ways have been suggested to quantify match importance. Goossens et al., (2012) take into account the number of teams that are still competing for a given prize. An integer program is used to determine the highest and lowest rank any team can still obtain at the end of the season when a certain match is played. Similar formulations are proposed by Ribeiro and Urrutia, (2005); Raack et al., (2014); Gotzes and Hoppmann, (2022). Geenens, (2014) measures the decisiveness of a match based on the expected entropy (i.e. uncertainty). Corona et al., (2017) use an entropy-based measure to identify the most decisive matches in the 2014 FIFA World Cup and the 2015 Copa América. Goller and Heiniger, (2023) extend the entropy measure by considering the difference between the reward probability distributions derived from the possible outcomes of a single event. Csató et al., 2024b introduce a classification scheme to determine weakly (where the final rank of one team is known, regardless of the match outcome) and strongly (where the match outcome does not influence the ranking of both teams) stakeless matches.
@@ 3 Tournament format
The comparison of tournament formats has received much attention in operational research since determining the format is probably the first task of the contest organizer. Even if the number of games that can be played is strictly limited, usually several options remain available. For example, Guyon, (2020) uncovers a serious drawback of the original format of the 2026 FIFA World Cup and lists 7 alternative options. Rennó-Costa, (2023) presents another one based on double elimination, while Guajardo and Krumer, (2023) develop three further formats. Choosing the best from such a set of feasible designs clearly requires their evaluation with respect to a number of criteria.
3.1 Efficacy
@@
At its inception, the tournament design literature was mainly concerned with the probability of the best player winning under a given tournament format, also called effectivity. Pioneering work was done by David, (1959), who used combinatorial techniques to compute exact winning probabilities for a SKO and SRR. To that end, a series of pairwise win probabilities needs to be assumed; $\pi_{ij}$ is the probability that $i$ beats $j$, which are collected in matrix $\Pi$. David, (1959) assumes a cumulative normal distribution $\Phi(\cdot)$ for the pairwise win probabilities, i.e. $\pi_{ij}=\Phi\{\frac{1}{\sigma}(r_{i}-r_{j})\}$, with $\sigma$ being the variance term and $r_{i}$ being the “true” rating of player $i$. Glenn, (1960) and Searls, (1963) apply a similar approach to evaluate tournament structures, although they assume an arbitrary matrix $\Pi$. All three papers compute the probabilities that other players finish in a certain position as well. Although limited to tournaments with 8 players, the authors find SKO to be the least effective format. Double and serial knockout formats are more effective, in particular, if matches are played as a best of three. For random knockout tournaments, Knuth and Lossers, (1987) compute the chance of player $i$ winning the tournament, assuming that it wins against player $i+1$ with probability $p$ and always defeats player $j$ if $j>i+1$. Adler et al., (2017) derive an upper and lower bound on the probability for of winning the tournament for all players, if player $i$ wins against player $j$ with probability $v_{i}/\left(v_{i}+v_{j}\right)$, where the values $v_{1}$, …, $v_{n}$ are exogenously given.
@@
As the exact computation of these expressions quickly becomes intractable, the standard approach in the literature is to perform Monte Carlo simulation. Using this approach, several papers conclude that round robin tournaments are typically more efficacious than knockout tournaments (Appleton, , 1995; McGarry and Schutz, , 1997; Ryvkin and Ortmann, , 2008; Scarf et al., , 2009; Ryvkin, , 2010; Scarf and Yusof, , 2011; Csató, 2021b, ). Furthermore, the efficacy of a SKO increases by extending it towards a DKO (McGarry and Schutz, , 1997; Stanton and Williams, , 2013). Efficacy can usually be improved by increasing the number of matches (Lasek and Gagolewski, , 2018) but the relationship is not necessarily monotonic for complex hybrid tournaments (Csató, 2021b, ). According to earlier studies (Appleton, , 1995; Dinh et al., , 2020), the Swiss system is rather weak in identifying the true ranking of players. However, Sziklai et al., (2022) find that the Swiss system outperforms other formats with the same number of games. To the best of our knowledge, the efficacy of incomplete round robin tournaments has not been examined yet.
@@
Table 1: Studies on the efficacy of tournament formats
(a) Papers focusing on efficacy
Study Format $\pi_{ij}$ Best Worst $n$
@@
SRR,
[…]
${}_{\text{S}}$, SRR+DKO,
${}_{\text{S}}$, DRR+DKO,
[…]
Scarf and Yusof, (2011) 23 formats Poisson()($\cdot$)( ⋅ ) SRR SKO 32
Stanton and Williams, (2013) SKO, DKO CRM, $\frac{1}{2}+\frac{j-i}{2(N-1)}$ DKO SKO 8–$10^{24}$
[…]
SRR, Swiss, Reaper T NS SRR SKO 8
[…]
D(8+6), D(4×$\times 7$),
D(4×$\times 6$)1 $\frac{(57-r_{i})^{k}}{(57-r_{i})^{k}+(57-r_{j})^{k}}$ D(8+6) D($4\times 6$) 24, 28
[…]
SRR, SKO, G64,
G66, G642 $\frac{1}{1+\left[\left(r_{i}+24\right)/\left(r_{j}+24\right)\right]}$ G66 G64 24
[…]
SKO-3, DKO, SRR+SKO $\frac{1}{1+10^{(r_{j}-r_{i})/400}}$ Swiss SKO 32
(b) Papers focusing on effectivity
@@
Unsurprisingly, there exist unavoidable trade-offs between these fairness criteria. The distance minimization and the break maximization problems are equivalent if the distance between every pair of teams is equal (Urrutia and Ribeiro, , 2006). Craig et al., (2009) focus on the trade-off between minimizing travel distance for teams, avoiding sequences of consecutive home or away games, and ensuring equity in rest times. For a SRR with $n$ teams, the minimal number of breaks is $n-2$, which is achieved (for instance) by the canonical schedule. However, this schedule also yields maximally unbalanced carry-over effects for an even number of teams (Lambrechts et al., , 2018). Günneç and Demir, (2019), Goossens et al., (2020), and Çavdaroğlu and Atan, (2022) discuss the trade-off between minimizing breaks and balancing the carry-over effect. Obviously, there are also trade-offs between fairness and organizational or financial criteria (see e.g. Goossens and Spieksma, (2009); Ribeiro and Urrutia, (2012)).
6.3 Attractiveness
In the sports scheduling literature, basically two approaches exist to deal with attractiveness. One approach assumes that the set of popular matches is known or given by a broadcaster, possibly inspired by studies like Wang et al., (2018); Krumer, (2020); Jakee et al., (2022). The schedule should then assign as many of these games as possible to commercially interesting timeslots, and avoid any overlap between them. This approach has been reported in several practical applications, from e.g. Australian football (Kyngäs et al., , 2017), football (Goossens and Spieksma, , 2009; Ribeiro and Urrutia, , 2012), and volleyball (Cocchi et al., , 2018). Matsumura and Hasuike, (2022) estimate the impact of scheduling decisions on TV audiences for the Japanese professional baseball league by multiple regression analysis to develop a schedule that maximizes the expected number of spectators. Domestic football competitions occasionally shut down for some days due to periods reserved for international matches played by national teams. According to Pérez, (2024), home wins have a lower probability after such an interruption, which leads to greater uncertainty of outcome. Consequently, the set of matches played immediately after “FIFA reserved dates” needs to be chosen strategically.
The other, much less attempted approach tries to find a schedule that is as attractive as possible given a prediction of match results, or robust with respect to a subset of possible outcomes. Yi, (2021) proposes an implementer-adversary approach to dynamically develop round robin schedules that maximize suspense, measured by how late in the season the winner is decided. The algorithm works by generating a set of scenarios (realizations of match outcomes) based on historical data from betting odds. The implementer problem finds a schedule that reveals the winner as late as possible, given the information about match outcomes. Next, the adversary problem finds a new (realistic) scenario under which the generated schedule is the worst possible, and this new scenario is added to the set of scenarios. The algorithm continues until the implementer and adversary solutions converge.
We see at least two avenues for future research on improving the attractiveness of tournament schedules. The first is to include match importance measures (see Section 2.3.3) in order to obtain a schedule that maximizes the total importance of matches over the season. This can be easily done in a SRR with three or four teams; papers focusing on stakeless matches usually recommend an optimal (Chater et al., , 2021; Guyon, , 2020) schedule. Using complete enumeration, Csató et al., 2024b have managed to solve this problem in a DRR with four teams and some scheduling constraints. Given the complexity of these measures, such an endeavor will be challenging computationally for larger tournaments. The second avenue focuses on dynamic scheduling, which means that the schedule is not computed and announced fully at the beginning of the season, but is released gradually during the season. Except for recent applications in the NFL (Bouzarth et al., , 2020), in the FIFA World Cup (Stronka, , 2024), and in e-sports (Dong et al., , 2023), dynamic scheduling has not received much attention in the literature. Yet, it may have a great potential to improve the attractiveness of tournaments, as recent information about the strength and ranking of players can be taken into account in order to maximize the importance of the remaining matches.
7 Ranking methods
The theoretical properties of ranking procedures have been widely discussed in the literature. González-Díaz et al., (2014) provide a comprehensive axiomatic analysis of ranking methods for general tournaments where the players may play an arbitrary number of matches against each other. Applying maximum likelihood to the (Zermelo–)Bradley–Terry model (Zermelo, , 1929; Bradley and Terry, , 1952), and the generalized row sum, a parametric family of ranking methods introduced by Chebotarev, (1994), turn out to be the most appealing by satisfying all reasonable axioms that are compatible with each other.
7.1 Efficacy
While efficacy in the context of tournament formats and seeding rules refers to the ability to rank the players according to their true (but hidden) strength, the efficacy of a ranking method usually refers to its predictive power, the ability of a ranking to accurately predict future matches. McHale and Davies, (2007) test a previous formula of the FIFA World Ranking from this perspective. Lasek et al., (2013) compare the predictive power of several methods for ranking national football teams. The Elo rating is found to be a better indicator of success than the old FIFA World Ranking, a result that is reinforced by Gásquez and Royuela, (2016). Indeed, FIFA reformed the World Ranking in 2018 (FIFA, , 2018). Nonetheless, the predictive power of the current FIFA World Ranking can still be considerably improved by incorporating home-field advantage and introducing a weighting by goal difference (Szczecinski and Roatis, , 2022). Somewhat analogously, a variant of the Elo rating (Football Club Elo Ratings, http://clubelo.com/) robustly outperforms the UEFA club coefficient in terms of predictive accuracy for the UEFA Champions League (Csató, 2024a, ).
Table 3: Papers on the efficacy of ranking methods
Tournament design is a research domain that has a strong link with practice. Practitioners and sports governing bodies will hopefully take an interest in this survey, too. In recent years, the results of academic research have frequently been adopted by sports administrators:
  • FIFA has introduced a better ranking procedure for men’s national teams in 2018 (FIFA, , 2018) (Section 2.2.3);
  • UEFA has modified the knockout bracket of the 2020 and 2024 UEFA European Championships to balance group advantages according to the recommendation of Guyon, 2018b (Section LABEL:Sec422);
  • FIFA has reformed the draw of the 2018 and 2022 FIFA World Cups inspired by Guyon, (2014), see Guyon, 2018a (Section LABEL:Sec52);
  • The proposal of Durán et al., (2017) for a fairer schedule of the South American Qualifiers for the 2018 FIFA World Cup has been unanimously approved by the participants and is currently being used (Section LABEL:Sec62).
Governing bodies in sports can also learn from each other. Both the FIBA Basketball World Cup and the FIFA World Cup aim to maximize geographic diversity in the group stage since maximizing intercontinental matches is useful both for attendance and ranking purposes (Guyon, , 2015). On the other hand, the 2023 World Men’s Handball Championship contained a group with two of the three African teams, another group with two of the four South American teams, while the number of European teams varied between one and three in the eight groups.
The example of the 2022 FIFA World Cup draw warns that it is worth consulting with the wider scientific community before rule changes are implemented. On 22 March 2022, FIFA released the details of this draw procedure, which treated the three placeholders of the playoffs unfairly and threatened the balancedness of the groups as discussed in Section LABEL:Sec51. Even though Csató, 2023b immediately realized the problem, it was clearly impossible to correct the mistake before the draw happened.
Consequently, since governing bodies in sports quite often change the design of their tournaments, usually driven by commercial motives, and some innovative ideas repeatedly emerge in practice, operational research in sports will certainly remain a growing research field for the coming decades.
Acknowledgements
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838, by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences, and by the Research Foundation - Flanders under Grant S005521N. We are grateful to András Gyimesi, Sergey Ilyin, and three anonymous reviewers for useful remarks.